NottheMonga2087:

Which expression shows the simplified form of (8r^5)^3 8r^15 8/r^15 512r^15 r^15/512

officialsavannah:

do the math, figure out all of the prombelms and which one is the same and simple as the expression

jhonyy9:

use this (x^a)^b = x^(a*b)

NottheMonga2087:

im confused

Angle:

\((r^5)^3 = ?\)

Angle:

do you know what that equals? ^

NottheMonga2087:

r ^15

Angle:

good now \(8^3 = ?\)

NottheMonga2087:

512

Angle:

\((8\times r^5)^3=(8)^3\times(r^5)^3 = \) combine the two answers you just got

Angle:

You ok? stuck? just look for the multiple choice option that has both of the answers from the previous comments

NottheMonga2087:

ok

NottheMonga2087:

d?

Angle:

close but not quite you want multiplication, because the 8 and the original r^5 were multiplied